OpenStudy (anonymous):

SOMEONE PLEASE HELP ME Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 (1 point) Question 2 options: 1) x = 0 2) x = 2 3) There is no x value as there is no solution. 4) x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (jdoe0001):

well.. if you solve the 1st equation "2x + y = 1" for "y", what would that give you?

OpenStudy (anonymous):

y=2x+1

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well.... hmmm how did you transfer the 2x over to the right side?

OpenStudy (anonymous):

ohhh I ment that as a negative sowwy y=-2x+1

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

woops got a bit caught up... yes is y = -2x+1 well then, now we know what "y" is equals to so \(\bf 2x + {\color{brown}{ y}} = 1\\ 4x + 2{\color{brown}{ y}} = -1 \\ \quad \\ \quad \\ 2x + {\color{brown}{ y}} = 1\implies {\color{brown}{ y}}=-2x+1\qquad thus \\ \quad \\ 4x + 2{\color{brown}{ y}} = -1\implies 4x + 2{\color{brown}{ (-2x+1)}} = -1\) so if you solve that for "x"... that'd give you what "x" is and then you can plug that in eihter equation, to solve for "y", to see what "y" is as a constant

OpenStudy (jdoe0001):

hmmm I get a funny result on that one

OpenStudy (jdoe0001):

\(\bf 4x + 2{\color{brown}{ y}} = -1\implies 4x + 2{\color{brown}{ (-2x+1)}} = -1\implies \cancel{ 4x-4x }+2=-1\) but... well obviously \(\large +2 \ne -1\) <--- so that's not really a valid EQUATion